A bat and ball together cost $31.00. The bat costs $15 more than the ball. What is the cost of one bat and two balls?


Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

A bat and ball together cost $31.00. The bat costs $15 more than the ball. What is the cost of one bat and two balls?
-------------
Bat + Ball = 31
Bat = Ball + 15
------
Ball+15 + Ball = 31
2*Ball = 16
The ball costs $8
etc

Answer by ikleyn(52754)

.

Write equations as you read the problem

    bat + ball = 31    (1)

    bat - ball = 15    (2)


Add the equations

    2*bat     = 31 + 15 = 46  ===>  bat = 46/2 = 23.


Then  from (1),  ball  = 31 - bat = 31 - 23 = 8.


Thus one bat costs $23;  one ball costs $8.


The rest is on you, since it is easy.

--------------

It can be solved in different ways, too;

but the way I showed in this post is the most educative (in my personal opinion).